OpenStudy (anonymous):

[In printing an article of 48,000 words, a printer decides to use two sizes of type. Using the larger type, a printed page contains 1,800 words. Using smaller type, a page contains 2,400 words. The article is allotted 21 full pages in a magazine. How many pages must be in smaller type?] / [Three individuals form a partnership and agree to divide the profits equally. X invests $9,000, Y invests $7,000, Z invests $4,000. If the profits are $4,800, how much less does X receive than if the profits were divided in proportion to the amount invested?] / {Can you clearly give me the details about above two questions, please? Thanks}

OpenStudy (anonymous):

Qns. 1 [In printing an article of 48,000 words, a printer decides to use two sizes of type. Using the larger type, a printed page contains 1,800 words. Using smaller type, a page contains 2,400 words. The article is allotted 21 full pages in a magazine. How many pages must be in smaller type?] Let x be the number of pages with the larger font size and y be the number of pages with the smaller font size. Form 2 equations relating both x and y to each other. x + y = 21 - (1) 1800 x + 2400y = 48000 - (2) From (1), (multiply by 1800) 1800 x + 1800 y = 37800 - (3) Take (2) subtract (3) 1800 x + 2400 y - 1800x - 1800 y = 48000 - 37800 600 y = 10200 y= 17 Substitute y = 17 to (1) x+ 17 = 21 x = 4 4 large pages and 21 smaller pages. Qns 2. [Three individuals form a partnership and agree to divide the profits equally. X invests $9,000, Y invests $7,000, Z invests $4,000. If the profits are $4,800, how much less does X receive than if the profits were divided in proportion to the amount invested?] X's profits when divided equally, 4800 /3 = 1600 X's profits when divided based on proportion invested, \[4800 \times \frac{9000}{9000+4000+ 7000}\] =2160. How much less does X receive? 2160 - 1600 = 560 dollars.